OpenStudy (candycanegone):

4/5 divided by 68 time 9

OpenStudy (sshayer):

let the speed of plane in calm atmosphere=x m/h speed of wind =y m/h 5x+5y=1050 x+y=210 7x-7y=1050 x-y=150 add the two and find x and then y

OpenStudy (candycanegone):

and i do that how?

OpenStudy (candycanegone):

12x +12y = 1050

OpenStudy (candycanegone):

?

OpenStudy (candycanegone):

@pooja195

OpenStudy (candycanegone):

@ganeshie8

OpenStudy (candycanegone):

help??

OpenStudy (candycanegone):

12x + 12y = 2100